OpenStudy (anonymous):

Solve these inequalities. (ii) 3^x +5 bigger or equal to 32 Please help! I don't know how to solve this though I a aware it involves logarithms. Please show me what to do and explain why.

Nnesha (nnesha):

\[\huge\rm 3^x +5 \ge 32\] yes right you have to take log both sides but first move `5` to the right side and then take log

Nnesha (nnesha):

make sense ?

OpenStudy (anonymous):

So It would be \[3^{x} \ge 17\] ? How do I take log of both sides?

Nnesha (nnesha):

17 ?? i guess thats a typo

OpenStudy (anonymous):

Um no actually. What did I do wrong?

Nnesha (nnesha):

how did you get 17 ?

OpenStudy (anonymous):

I subtracted 5 from both sides?

Nnesha (nnesha):

yes right so 32-5 isn't equal to 17 :=)

OpenStudy (anonymous):

Ooooh. I do mistakes like that often. Oops!

OpenStudy (anonymous):

ok so after \[3^{x} \ge 27\] how do I take log of both sides?

Nnesha (nnesha):

yes right now take log both sides and apply the product rule \[\huge\rm log (3)^x \ge \log(27)\]

Nnesha (nnesha):

power rule ***

OpenStudy (anonymous):

Ok I am new too logarithms so sorry for taking so much time. If you use the power rule does that mean that I move the exponent x in front of the first logarithm?

Nnesha (nnesha):

yes right :=) and its okay take ur time!

OpenStudy (anonymous):

Thanks! All right so, \[x \log_{}3 \ge \log27\]

Nnesha (nnesha):

yes right now to solve for x you should move log(3) to the right side :=) and then use calculator :P that's it!

OpenStudy (anonymous):

Ok, thank you so much! :D

Nnesha (nnesha):

what did you get ?? o.O

OpenStudy (anonymous):

x\[x \ge 3\]

Nnesha (nnesha):

yes right!

Nnesha (nnesha):

good job!

OpenStudy (anonymous):

Hooray!